OpenStudy (anonymous):

What is the quadratic function that is created with roots at 2 and 4 and a vertex at (3, 1)?

OpenStudy (anonymous):

\[f(x)=a(x-2)(x-4)\] and solve for a via \[f(3)=1=a(1-2)(1-3)\]

OpenStudy (anonymous):

gives \[1=2a\] or \[a=\frac{1}{2}\]

OpenStudy (anonymous):

you could have also written \[f(x)=a(x-3)^2+1\] and solved for "a" by writing \[f(2)=0=a(2-3)^2+1\] hold the phone, i get a different "a"

OpenStudy (anonymous):

are you sure this question is written correctly??

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

it is...

OpenStudy (anonymous):

maybe my algebra is messed up, let me try again

OpenStudy (wasiqss):

satellit plz try my question again

OpenStudy (anonymous):

go away! he's mine!

OpenStudy (wasiqss):

no he is mine :P http://openstudy.com/study#/updates/4f0c8288e4b084a815fc61e9

OpenStudy (anonymous):

Thieves!

OpenStudy (anonymous):

oh yeah my algebra was messed up sorry

OpenStudy (anonymous):

\[f(3)=1=a(3-2)(3-4)=1\] so \[-a=1\] ir \[a=-1\] just like with \[f(2)=0=a(2-3)^2+1\] \[0=a+1\] \[a=-1\] sorry

OpenStudy (anonymous):

so answer is \[f(x)=-(x-2)(x-4)=-(x-3)^2+1=-x^2+6x-8\]

OpenStudy (anonymous):

Awesome! Thank you for your help and explaining everything! Long time no see man!